Since the 1980s, HIV (Human Immunodeficiency Virus) has been infecting humans around the world causing the condition known as AIDS (Acquired Immune Deficiency Syndrome). HIV, like all viruses, needs to enter cells and use their machinery to reproduce and spread. During HIV infection, the virus enters specific cells of the immune system (T-cells) by "docking" onto cell surface proteins, including one called CCR5 Genetic analysis of individuals who are naturally immune (resistant) to HIV have revealed that resistance to HIV is due to a mutation in the gene coding for the CCR5 protein. The mutant allele is called ccr5-delta32 (432, for short). The A32 mutation is a deletion of 32 nucleotides in the middle of the CCR5 gene sequence, resulting in a defective protein that does not localize to the surface of T-cells. The A32 allele is recessive to the wild-type allele; individuals who are homozygous for the A32 allele are resistant to HIV infection Barring the Door | How a mutation can protect against HIV HIV Without CCRS With CCR5 To enter cells, one Some people are born with a mutation that thing that most ust do is bind to CCR5, a molecule appearing on their cells, making them reabt mental White hlood cell from the calls surface. gene therapy approaches aim to mimic this mutation in AIDS patients. Centor Note: illustration is schematic The table below is an excerpt of a paper providing the results of genetic analysis of population samples in European countries in 2002. Country Town Genotypes N IMPORTANT NOTE: the table reports the sample size (N) and the number of heterozygous (+/A32) and homozygous mutant (A32/A32) individuals for each population sampled. A plus sign (+) is used in the table to signify the wild-type allele, which we typically label as wt. (region/group) +432 4321432 Iceland 102 24 204 46 6 Sweden Umea Norway Finland Oslo 100 21 0 Helsinki 98 31 Lithuania Vilnius 283 61 2 Ireland 44 4 0 283 57 3 Britain Denmark Copenhagen Moscow Brussels 239 49 5 Adapted from: Lucotte (2002), Infection, Genetics and Evolution 1: 201-205. Russia 176 43 8 Belgium 704 114
Model Rubric This rubric will be used to evaluate the model you build in response to question 4. Refer to the rubric to evaluate your work before you turn in the assignment. Model elements Expectations (1 point each; total = 10 pts) (Abbreviations) 1 The model clearly shows what the genotype of the organism is (both alleles) 2 All required structures (allele, gene, chromosome) are included. 3 - The relationships between alleles, gene and chromosome are accurate. Genotype (Gen) 1 The model clearly shows transcription of each allele. 2 The mechanisms/processes illustrating transcription are accurate. Transcription (Txn) 1 The model clearly shows how translation results in specific proteins. 2 The mechanisms/processes illustrating translation are accurate. Translation (TIn) 1 The model clearly shows what the phenotype of the organism is. 2 The phenotype is illustrated as a result of the protein function 3-The mechanisms/processes illustrating the protein function are accurate. Phenotype (Phe)
[2 pts] What were the frequencies of the mutant A32 allele and the wild-type CCR5 (wf) allele in the population sample from Sweden in 2002? Show your work. 1. 2. [6 pts] Was the Sweden population in Hardy-Weinberg Equilibrium for the CCR5 locus in 2002? Make sure you clearly state "yes" or "no", show ALL necessary evidence and explain your reasoning, for full credit. [The critical X2 value for this analysis is X2=3.84, p = 0.05]. wt/wt wt/A32 432/132 Observed Expected (O-E)/E X2 = 3. Suppose that the population of Sweden is sampled again, this year (2019), and that the frequency of the 432 allele has not significantly changed since the time the initial paper was published (2002) Based on this information alone, you can confidently claim that: a. [3 pts] the Sweden population has evolved for the CCR5 locus. OTrue False Briefly explain your reasoning. b. [3 pts] the Sweden population is in Hardy-Weinberg equilibrium in 2019. OTrue False Briefly explain your reasoning.
4. [10 pts] Individuals who are heterozygous for the A32 are susceptible to HIV infection. In the space below, draw a model representing HOW genotype determines phenotype in individuals who are heterozygous for the 432 allele. Your model must include, at a minimum, the following structures: chromosome, gene, allele, mRNA, protein. Modify these terms as necessary to create a model that is specific to the function of CCR5 in humans. To create a meaningful model of a biological system, you will need to: 1. represent all structures in boxes and mechanisms (or relationships) on arrows; 2. connect all boxes together within a single model; 3. make sure each arrow has a direction (meaning it is not just a line!); 4. be able to read each "box-arrow-box" group as a complete, coherent statement; and 5. build the model so that it illustrates the function of the system.
Answer #1

Sweden population n= 204.

Double mutants = 6

Heterozygote = 46

Homozygotes wild = 204-(6+46)= 152

Each mutant has two mutant alleles and each heterozygote has one mutant alleles.

So frequency of mutant allele= 2*6+1*46= 12+46=58

Total alleles = 2*204=408

Frequency = 58/408= 0.142.

Frequency of wild type allele = 2*152+1*46=304+46=350

Frequency of wild allele = 350/408= 0.857

  1. Herdey Weinberg equilibrium says that sum of all frequencies is equal to 1. So wild allele + mutant allele frequency = 1.

0.142+0.857= 0.999

Yes, the Sweden population is HW equilibrium.

Solution:

Observed frequencies = frequency of mutant = q^2 = 6/204= 0.029.

q= root of 0.029

q= 0.17

So, p = frequency of wild allele = 1-0.17=0.83

So gene frequency of wild p^2= 0.83^2= 0.689

The gene frequency if heterozygote = 2pq= 2*0.83*0.17= 0.282

Expected is (from question 1), p= 0.857, p^2= 0.734

q= 0.142, q^2= 0.02

Heterozygote 2pq= 2(0.857)(0.142)=0.243.

Now, doing chi square analysis:

(O-E)^2/E

= [(0.689-0.734)^2/0.734] + [(0.029-0.02)^2/0.02] + [(0.282-0.243)^2/0.243]

= 0.002+0.004+0.006

= 0.012 this is the X2 value.

As we see that this value is less than 3.84. That means there is no significant difference between the two values. In other words, Sweden population follow HW equilibrium.

  1. In the new population, ten frequency of mutant allele has not changed significantly. That means there is no selection in the favour of this allele. It also means that wild type allele is not favored wither, otherwise the mutant allele frequency would have gone down. So the Answer is false.

The population continues to to be HW equilibrium, since the frequencies have not changed.

  1. Please find the model attached below.

Helerogygos f D32 432/CCRS CA/) + Trenniphen Doemit allos Binding Amtant prtein MRNA toild CCRS Alloces binding prlen HIV No